# Orient Anything

### v1

@@ 3.2 Implementation
Informed by our insights from Section 3.1, we now present our state-of-the-art method for orientation estimation. Our pipeline consists of two components. Our first component, which we call the quotient orienter, is a neural network trained to solve Problem 3. We quotient the $L^{2}$ objective by ^$\hat{\mathcal{R}}:=\mathcal{O}\subseteq SO(3)$, the octahedral group containing the 24 rotational symmetries of a cube. This is among the largest finite subgroups of $SO(3)$ (only the cyclic group $C_{n}$ for $n\geq 48$ and dihedral group $D_{n}$ for $n\geq 4$ can contain more subgroups), and it includes many rotational symmetries that commonly occur in real-world shapes. Our second component, which we call the flipper, is a neural network trained to solve the classification problem 5. We illustrate the output of each stage of this pipeline in Figure LABEL:fig:our-pipeline-bench. As many shapes possess multiple plausible orientations, we use conformal prediction to enable our flipper to output adaptive prediction sets whose size varies with the flipper model’s uncertainty. We provide further implementation details below.
@@
We also parametrize our flipper by a DGCNN operating on point clouds. We train the flipper on point clouds sampled from the surface of Shapenet meshes by optimizing Equation 5. We draw rotations $Q\sim U(\mathcal{O})$ during training, and simulate inaccuracies in our quotient orienter’s predictions by further rotating the training shapes about a randomly drawn axis by an angle uniformly drawn from $[0,10]$ degrees. We provide full architecture and training details in Appendix B.
@@
To mitigate this issue in pipelines with a human in the loop, we enable our flipper model to output adaptive prediction sets whose size varies with the flipper’s uncertainty (Romano et al., 2020). This method uses a small conformal calibration set drawn from the validation data to learn a tuning parameter $\tau>0$ that controls the size of the prediction sets. Given the flipper model’s output probabilities $p_{\phi}(S)\in\Delta^{|\hat{\mathcal{R}}|-1}$ for some shape $S$, one sorts $p_{\phi}(S)$ in descending order and adds elements of $\hat{\mathcal{R}}$ to the prediction set until their total mass in $p_{\phi}(S)$ reaches $\tau$. Intuitively, these sets will be small when the flipper is confident in its prediction and assigns large mass to the highest-probability classes. Conversely, the sets will be large when the flipper is uncertain and assigns similar mass to most classes.
@@ 4 Experiments
We now evaluate our method’s performance on orientation estimation. We first follow the evaluation procedure in Pang et al. (2022) and benchmark against their “Upright-Net,” which represents the current state of the art for orientation estimation. Upright-Net can only map shapes into upright orientation, where a shape’s up axis is aligned with the $y$-axis; in contrast, our method recovers a full orientation $\Omega_{S}$ for each shape. We therefore follow this benchmark with an evaluation of our method on the more challenging task of full-orientation estimation. We incorporate adaptive prediction sets at this stage and demonstrate that our method reliably provides a plausible set of candidate orientations for diverse shapes unseen during training.
@@ 4.1 Upright orientation estimation
We then follow the evaluation procedure in Pang et al. (2022) to benchmark our method against their SOTA method for upright orientation estimation. We randomly rotate shapes $S$ in the validation set, use our two-stage pipeline and Upright-Net to estimate the up-axis $\omega_{RS}^{y}$ of each randomly rotated shape $RS$, and then measure the angular error $\textrm{arccos}(\langle\hat{\omega}_{RS}^{y},\omega_{RS}^{y}\rangle)$ between the estimated and ground truth up-axis. Our method’s estimated up-axis is the second column of our estimated orientation matrix $\hat{\Omega}_{RS}$. In contrast, Upright-Net predicts a set of base points for $RS$, fits a plane to these points, and returns this plane’s normal vector pointing towards the shape’s center of mass. We follow Pang et al. (2022) and define our methods’ respective accuracies to be the proportion of validation meshes whose angular error is less than $10^{\circ}$.
@@
Upright-Net 69.5 % 62.3 %
We depict the results of this benchmark in Table 1. Our method improves on Upright-Net’s up-axis estimation accuracy by nearly 20 percentage points, corresponding to a 64.6% reduction in the error rate relative to the previous state of the art. To provide a more comprehensive picture of our respective models’ performance, we also report angular loss histograms for our model and Upright-Net in the left panel of Figure LABEL:fig:up-axis-loss-histograms. Our model primarily fails by outputting orientations that are $90^{\circ}$ or $180^{\circ}$ away from the correct orientation, which correspond to failures of the flipper. In contrast, Upright-Net’s failures are more evenly distributed across angular errors. Finally, we depict a grid of non-cherry-picked outputs of our model and Upright-Net in Figure LABEL:fig:sec-4.1-grid and highlight each model’s failure cases in red.
@@
These results demonstrate that our method significantly improves over the state of the art in upright orientation estimation. In the following section, we show that our method also successfully recovers the full orientation $\Omega_{RS}$ of a rotated shape, a more challenging task than the well-studied task of upright orientation estimation. Using our estimated orientations, we return a wide variety of shapes into canonical orientation.
@@ 4.2 Full-orientation estimation
We now evaluate our method’s performance on full-orientation estimation, in which we use our model’s full orientation matrix $\hat{\Omega}_{RS}$ to transform an arbitrarily-rotated shape $RS$ to the canonically-oriented shape $S$. To our knowledge, our algorithm is the first to solve this task for generic shapes without requiring class information at training time or at inference time. We now record the angular distance between our estimated orientations $\hat{\Omega}_{RS}$ and the ground truth orientations $\Omega_{RS}=R\Omega_{S}=R$. This angular distance is defined as $d(\hat{\Omega}_{RS},R):=\textrm{arccos}(\frac{\textrm{tr}R_{\textrm{diff}}-1}{2})$, where $R_{\textrm{diff}}:=\hat{\Omega}_{RS}R^{\top}$.
As noted in Section 3.2, many real-world shapes have several plausible canonical orientations, and our flipper may also map nearly-symmetric shapes to a uniform distribution over their near-symmetries. In particular, while most real-world shapes have a well-defined upright orientation, their front-facing orientation is often ambiguous. To account for this, we incorporate adaptive prediction sets at this stage of our evaluation. We measure the angular distance between the estimated orientations $\hat{\Omega}^{k}_{RS}$ corresponding to the top $K=4$ flips in our flipper’s model distribution and the ground truth orientation $\Omega_{RS}$, and take their maximum to obtain our reported angular errors. This resolves ambiguities in a shape’s front-facing direction (there are 4 possible front-facing directions for each upright orientation) and also simulates the ability of a human in the loop to choose between a small set of candidate orientations.
We describe our method’s performance on full-orientation estimation in Table 2, where we adopt a $10^{\circ}$ accuracy threshold for full rotation estimation. Our method achieves high top-4 accuracy on full-orientation estimation, but its accuracy deteriorates in the $K=1$ case, where one can only consider the flipper’s highest-probability class. This is partially attributable to ambiguities in the front-facing orientation of many shapes. To demonstrate this, the left panel of Figure LABEL:fig:sec-4.2-grid depicts shapes where our method’s top-1 angular error is $>10^{\circ}$ in red. These failures primarily correspond to shapes that are symmetric with respect to rotations about the $y$-axis; these shapes are correctly oriented even though our method has recovered incorrect orientations.
Table 2: Full-orientation estimation accuracy
Method Accuracy ($\uparrow$)
Top-1 Top-4
Ours 68.3 % 85.8 %
Figure LABEL:fig:sec-4.2-grid compares the shapes obtained using our model’s outputs when $K=1$ (where we only consider the flipper’s highest-probability predicted flip) and when $K=4$ (where we depict the best flip among the top 4 classes in the flipper’s model distribution) to the canonically-oriented shapes. We highlight the model’s failure cases in red. Even in the $K=1$ case, many of the model’s failures correspond to orientations that are plausible or correct up to a symmetry of the shape. We bolster this claim with additional non-cherry-picked examples in Appendix C; see Figures 10, 11.
@@
Finally, in Figure 12 in Appendix C, we depict transformed shapes obtained by applying our method to randomly-rotated shapes from the Objaverse dataset (Deitke et al., 2023). This dataset contains highly diverse meshes of varying quality and therefore serves as a useful test case for our method’s performance on out-of-distribution shapes. (As these meshes are not consistently oriented in the dataset, we cannot train on them or report meaningful error metrics.) Using our orientation pipeline, one reliably recovers shapes that are canonically-oriented up to an octahedral flip. Our flipper has greater difficulty handling out-of-distribution meshes, but predicts an acceptable flip in many cases. We expect that training our flipper on a larger dataset of oriented shapes will further improve its generalization performance.
@@ 5 Conclusion
This work introduces a state-of-the-art method for 3D orientation estimation. Whereas previous approaches can only infer upright orientations for limited classes of shapes, our method successfully recovers entire orientations for general shapes. We show that naïve regression-based approaches for orientation estimation degenerate on rotationally-symmetric shapes, which are common in practice, and develop a two-stage orientation pipeline that avoids these obstacles. Our pipeline first orients an arbitrarily rotated input shape up to an octahedral symmetry, and then predicts the octahedral symmetry that maps the first-stage output to the canonically-oriented shape.
In the language of group theory, our two-stage approach first learns to map an input shape $RS$ onto the orbit of the desired output $S$ under some symmetry group $\hat{\mathcal{R}}$, and then learns a distribution over elements of this orbit. We anticipate that this factorization of geometric learning problems will be broadly applicable throughout 3D deep learning for tackling problems that are ill-posed due to the presence of symmetries. We also believe that our results can be further improved by training our quotient orienter and flipper models on larger datasets of consistently-oriented shapes as they become available.
Acknowledgments
We thank the authors of Pang et al. (2022) for their gracious assistance with their source code. CS thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship.
Appendix A Proofs
A.1 Proposition 3.1
Suppose $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$, and fix some non-identity rotation $R\in\mathcal{R}_{S}$ under which $S$ is symmetric. Then $RS=S$, but $f(RS)=R\Omega_{S}\neq\Omega_{S}=f(S)$, so $f(RS)\neq f(S)$ even though $RS=S$. Hence $f$ must be a one-to-many map and is therefore not a function. $\blacksquare$
@@ Appendix A Proofs
Hence $L^{2}$ regression fails to learn the orientation $\Omega_{S}$ of a shape $S\in\mathcal{S}$ that possesses a non-trivial set of rotational symmetries $\mathcal{R}_{S}$. $\blacksquare$
A.3 Proposition 3.3
@@ B.3 Adaptive prediction sets
Appendix C Additional results
Figure 10: A grid of our model’s top-1 outputs given randomly rotated, non-cherrypicked meshes from Shapenet. We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view and highlight failure cases in red. Even when our model fails to recover the correct orientation, the recovered shape is often acceptable in practice.
Figure 11: A grid of our model’s top-4 outputs given randomly rotated, non-cherrypicked meshes from Shapenet. We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view and highlight failure cases in red.
Figure 12: A grid of our model’s top-1 outputs given randomly rotated, non-cherrypicked meshes from Objaverse. We depict a front view of our model’s recovered shapes along with an inset depicting an isometric view. Our quotient regressor consistently succeeds on out of distribution meshes, as most of our pipeline’s outputs are correctly oriented up to a cube flip. Our flipper has greater difficulty generalizing, but predicts an acceptable flip in many cases. We expect that training on a larger and more diverse dataset of oriented shapes will improve our flipper’s generalization performance.

### v2

@@ 3.4 Implementation
Informed by our insights from Sections 3.2 and 3.3, we now present our state-of-the-art method for orientation estimation. Our pipeline consists of two components. Our first component, which we call the quotient orienter, is a neural network trained to solve Problem 3. We quotient the $L_{2}$ objective by ^$\hat{\mathcal{R}}:=\mathcal{O}\subseteq SO(3)$, the octahedral group containing the 24 rotational symmetries of a cube. This is among the largest finite subgroups of $SO(3)$ (only the cyclic group $C_{n}$ for $n\geq 48$ and dihedral group $D_{n}$ for $n\geq 4$ can contain more subgroups), and it includes many rotational symmetries that commonly occur in real-world shapes.
Our second component, which we call the flipper, is a classifier trained to predict the rotation $Q^{*}\in\hat{\mathcal{R}}$ from the output $f^{*}(RS)=RQ^{*}$ of the quotient regression model. We illustrate the output of each stage of this pipeline in Figure LABEL:fig:our-pipeline-bench. As many shapes possess multiple plausible orientations, we use conformal prediction to enable our flipper to output adaptive prediction sets whose size varies with the flipper model’s uncertainty. We provide further implementation details below.
@@
We also parametrize our flipper by a DGCNN operating on point clouds. We train the flipper on point clouds sampled from the surface of Shapenet meshes by solving the $|\hat{\mathcal{R}}|$-class classification problem described in Section 3.3. We draw rotations $Q\sim U(\mathcal{O})$ during training, and simulate inaccuracies in our quotient orienter’s predictions by further rotating the training shapes about a randomly drawn axis by an angle uniformly drawn from $[0,10]$ degrees. We provide full architecture and training details in Appendix B.
@@
To mitigate this issue in pipelines with a human in the loop, we enable our flipper model to output adaptive prediction sets whose size varies with the flipper’s uncertainty (Romano et al., 2020). This method uses a small conformal calibration set drawn from the validation data to learn a threshold parameter $\tau>0$ associated with a coverage probability $\alpha\in(0,1)$ that controls the size of the prediction sets. Given the flipper model’s output probabilities $p_{\phi}(S)\in\Delta^{|\hat{\mathcal{R}}|-1}$ for some shape $S$, one sorts $p_{\phi}(S)$ in descending order and adds elements of $\hat{\mathcal{R}}$ to the prediction set until their total mass in $p_{\phi}(S)$ reaches $\tau$. Intuitively, these sets will be small when the flipper is confident in its prediction and assigns large mass to the highest-probability classes. Conversely, the sets will be large when the flipper is uncertain and assigns similar mass to most classes.
@@ 4 Experiments
We now evaluate our method’s performance on orientation estimation. We first follow the evaluation procedure in Pang et al. (2022) and benchmark against their “Upright-Net,” which represents the current state of the art for orientation estimation. Upright-Net can only map shapes into upright orientation, where a shape’s up-axis is aligned with the $y$-axis; in contrast, our method recovers a full orientation $\Omega_{S}$ for each shape. We therefore follow this benchmark with an evaluation of our method on the more challenging task of full-orientation estimation. We incorporate adaptive prediction sets at this stage and demonstrate that our method reliably provides a plausible set of candidate orientations for diverse shapes unseen during training. We train and evaluate all models on Shapenet (Chang et al., 2015), as this is the largest and most diverse dataset we are aware of containing canonically oriented shapes. However, we report qualitative results for our method’s out-of-distribution performance on Objaverse in Appendix C.
@@ 4.1 Upright orientation estimation
Upright-Net (Pang et al., 2022) 69.5 % 62.3 %
We then follow the evaluation procedure in Pang et al. (2022) to benchmark our method against their SOTA method for upright orientation estimation. We randomly rotate shapes $S$ in the validation set, use our two-stage pipeline and Upright-Net to estimate the up-axis $\omega_{RS}^{y}$ of each randomly rotated shape $RS$, and then measure the angular error $\textrm{arccos}(\langle\hat{\omega}_{RS}^{y},\omega_{RS}^{y}\rangle)$ between the estimated and ground truth up-axis. Our method’s estimated up-axis is the second column of our estimated orientation matrix $\hat{\Omega}_{RS}$.
In contrast, Upright-Net predicts a set of base points for $RS$, fits a plane to these points, and returns this plane’s normal vector pointing towards the shape’s center of mass. This method relies on a restrictive prior on the geometry of the input shapes that does not hold for shapes which do not naturally lie on a supporting base. We follow Pang et al. (2022) and define our methods’ respective accuracies to be the proportion of validation meshes whose angular error is less than $10^{\circ}$.
We depict the results of this benchmark in Table 1. Our method improves on Upright-Net’s up-axis estimation accuracy by nearly 20 percentage points, corresponding to a 64.6% reduction in the error rate relative to the previous state of the art. To provide a more comprehensive picture of our respective models’ performance, we also report the empirical CDF of angular errors for our model and Upright-Net in the left panel of Figure LABEL:fig:up-axis-loss-histograms. Our model primarily fails by outputting orientations that are $90^{\circ}$ or $180^{\circ}$ away from the correct orientation, which correspond to failures of the flipper. In contrast, Upright-Net’s failures are more evenly distributed across angular errors. Finally, we depict a grid of non-cherry-picked outputs of our model and Upright-Net in Figure LABEL:fig:sec-4.1-grid and highlight failure cases in red.
@@
These results demonstrate that our method significantly improves over the state of the art in up-axis estimation. In the following section, we show that our method also successfully recovers the full orientation $\Omega_{RS}$ of a rotated shape up to its symmetries, a more challenging task than upright orientation estimation. Using our estimated orientations, we return a wide variety of shapes into canonical orientation.
@@ 4.2 Full-orientation estimation
We now evaluate our method’s performance on full-orientation estimation, in which we use our model’s full predicted orientation matrix $\hat{\Omega}_{RS}$ to transform an arbitrarily-rotated shape $RS$ to the canonically oriented shape $S$. We draw shapes $S$ from the Shapenet validation set and randomly rotate them to obtain inputs $RS$ to our pipeline. Because several orientation matrices may be associated with a single canonically oriented shape when this shape has rotational symmetries, we report the symmetric chamfer distance between our pipeline’s output shape and the ground truth shape. Because this is a shape-to-shape metric, it is insensitive to rotational symmetries: If $RS=S$, then $d(S,T)=d(RS,T)$, where $d$ is the chamfer distance and $T$ is a reference shape.
To our knowledge, our algorithm is the first to solve this task for generic shapes without requiring class information at training time or at inference time. We therefore construct two baselines using Upright-Net. As this algorithm only predicts a shape’s up-axis (the second column of its orientation matrix), we must augment Upright-Net’s predictions with a front-axis and a side-axis. We generate a baseline called “Upright-Net-Oracle” by augmenting Upright-Net’s predicted up-axis with the shape’s ground truth front-axis, and a second baseline called “Upright-Net-Random” by augmenting the predicted up-axis with a random vector that we orthogonalize with respect to the up-axis. In both cases, we obtain the side-axis by taking the cross product of the respective up- and front-axes. “Upright-Net-Oracle” upper-bounds the performance of a hypothetical orienter built on Upright-Net, whereas “Upright-Net-Random” provides a lower bound on such an orienter’s performance.
We benchmark our pipeline’s performance (“Orient Anything”) against each baseline in Table 2, where we report the mean chamfer distance between each model’s predicted shapes and the corresponding ground truth shapes, which are in canonical orientation. Orient Anything achieves an 84% reduction in mean chamfer distance between predicted and ground truth shapes relative to the oracle baseline, which augments Upright-Net’s predicted up-axis with the shape’s ground truth front-axis. In practice, one does not have access to this front-axis at inference time, so this represents a loose upper bound on the potential performance of a shape orientation algorithm built upon Upright-Net’s predicted up-axes.
This empirically validates our Proposition 3.2 and demonstrates that our pipeline successfully recovers canonically oriented shapes, even when symmetries preclude the orienter map from being represented by a function. We further illustrate our pipeline’s performance in Figure 9 in Appendix C, where we depict a large grid of shapes that have been oriented by our pipeline. Even when our model’s outputs disagree with the ground truth pose in Shapenet (we render the corresponding meshes in red), the recovered shape is often acceptable in practice.
@@
Finally, in Figure 11 in Appendix C, we depict transformed shapes obtained by applying our method to randomly-rotated shapes from the Objaverse dataset (Deitke et al., 2023). This dataset contains highly diverse meshes of varying quality and therefore serves as a useful test case for our method’s performance on out-of-distribution shapes. (As these shapes are not canonically oriented, we cannot train on them or report meaningful error metrics.) Using our orientation pipeline, one reliably recovers shapes that are canonically oriented up to an octahedral flip. Our flipper has greater difficulty handling out-of-distribution meshes, but predicts an acceptable flip in many cases. We expect that training our flipper on a larger dataset of oriented shapes will further improve its generalization performance.
Table 2: Full-orientation estimation performance on the Shapenet validation set.
Method Mean chamfer distance ($\downarrow$)
Upright-Net-Random 0.10801 $\pm$ 0.13824
Upright-Net-Oracle 0.05481 $\pm$ 0.13016
Orient Anything 0.00856 $\pm$ 0.0396
Orient Anything-APS 0.00208 $\pm$ 0.01407
Figure 8: Histogram of APS sizes with coverage probability $\alpha=0.3$. The median APS contains 2 shapes, and 90% of APSs contain at most 8 shapes.
Adaptive prediction sets.
Many real-world shapes have several plausible canonical orientations, even if they lack rotational symmetries. In particular, while most real-world shapes have a well-defined upright orientation, their front-facing orientation is often ambiguous. In Section 3.4, we proposed enabling a human in the loop to resolve these ambiguities by having our orientation pipeline output adaptive prediction sets (APSs), whose size varies with the flipper’s uncertainty. We now incorporate this technique into our evaluation. We learn a threshold parameter $\tau$ corresponding to a coverage probability of $\alpha=0.3$, and at inference time output a set of flips in descending order of probability under the flipper’s output distribution until their total probability mass reaches $\tau$. Intuitively, this set will be small when the flipper is confident in its prediction and large otherwise.
By applying each flip to the first-stage orienter’s output, we obtain a shape orientation pipeline that outputs sets of candidate shapes $\hat{S}_{c}$ given a single arbitrarily-rotated input $RS$. We call this method “Orient Anything-APS.” To evaluate its performance, we compute the chamfer distance between each candidate shape and the ground truth shape, and take the minimum over these chamfer distances. As one may trivially reduce the minimum chamfer distance over an APS by outputting arbitrarily large sets, we report a histogram of APS sizes in Figure 8 to demonstrate that these sets are typically small.
We report the performance of “Orient Anything-APS” in the final row of Table 2. Allowing our pipeline to output an APS reduces the mean chamfer distance between the closest shape in this set to the ground truth by a factor of 4 relative to Orient Anything without this feature. Figure 8 shows that with our chosen coverage probability, most APSs are small; the median APS size is 2, and 90% of APSs have at most 8 shapes. This demonstrates that our flipper typically assigns most of its probability mass to a few octahedral symmetries, which often contain a rotation that returns the shape to its canonical pose in Shapenet. In Figure 10 in Appendix C, we also include a grid of meshes in the prediction sets associated with 20 randomly-selected validation shapes. These prediction sets typically consist of a small collection of plausible shape orientations; by manually inspecting them, a human in the loop may select the orientation that is most suitable for their target application.
@@ 5 Conclusion
This work introduces a state-of-the-art method for 3D orientation estimation. Whereas previous approaches can only infer upright orientations for limited classes of shapes, our method successfully recovers entire orientations for general shapes. We show that naïve regression-based approaches for orientation estimation degenerate on rotationally-symmetric shapes, which are common in practice, and develop a two-stage orientation pipeline that avoids these obstacles. Our pipeline first orients an arbitrarily rotated input shape up to an octahedral symmetry, and then predicts the octahedral symmetry that maps the first-stage output to the canonically oriented shape. We anticipate that this factorization of geometric learning problems will be broadly applicable throughout 3D deep learning for tackling problems that are ill-posed due to the presence of symmetries.
Impact statement.
This paper presents a novel algorithm for 3D shape orientation. We do not believe that there are notable ethical considerations or societal impacts of our work meriting specific consideration.
Acknowledgments.
We thank the authors of Pang et al. (2022) for their gracious assistance with their source code and Justin Solomon for valuable discussions. CS thanks Backflip AI for hosting him and generously supporting him with mentorship and compute throughout his internship.
Appendix A Additional results and proofs
A.1 The orienter map for symmetric shapes is not a function
Proposition A.1.
Proof: Suppose $f(RS)=R\Omega_{S}$ for all $R\in SO(3)$, and fix some non-identity rotation $R\in\mathcal{R}_{S}$ under which $S$ is symmetric. Then $RS=S$, but $f(RS)=R\Omega_{S}\neq\Omega_{S}=f(S)$, so $f(RS)\neq f(S)$ even though $RS=S$. Hence $f$ must be a one-to-many map and is therefore not a function. $\blacksquare$
@@ A.1 The orienter map for symmetric shapes is not a function
Hence $L_{2}$ regression fails to learn the orientation $\Omega_{S}$ of a shape $S\in\mathcal{S}$ that possesses a non-trivial set of rotational symmetries $\mathcal{R}_{S}$. $\blacksquare$
A.3 Non-uniqueness of solution to Problem 2
Consider the bench shape $S$ depicted in Figures LABEL:fig:teaser, 2, LABEL:fig:prop3.1. As shown in Figure LABEL:fig:prop3.1, this shape has two rotational symmetries: The identity rotation, and a $180^{\circ}$ rotation about the $y$-axis. One may represent these rotations by the matrices $I$ and $Q:=(-e_{x},e_{y},-e_{z})$, respectively, where $e_{x},e_{y},e_{z}$ are the standard basis vectors.
Proposition 3.1 states that one solves the $L^{2}$ regression problem 2 for the bench shape by computing the arithmetic mean of $I$ and $Q$ and then orthogonally projecting this matrix onto $SO(3)$. The arithmetic mean of $I,Q$ is the matrix $M:=(0,e_{y},0)$, and one computes its orthogonal projection onto $SO(3)$ by solving a special Procrustes problem (Gower & Dijksterhuis, 2004):
$\min_{R\in SO(3)}\|R-M\|_{F}^{2}$ (6)
The solution to this problem is non-unique for $M:=(0,e_{y},0)$, and the minimum is attained by any rotation about the $y$-axis (i.e. any rotation matrix whose second solumn is $e_{y}$). This shows that even a single non-trivial rotational symmetry leads to an entire submanifold of solutions $f^{*}(S)$ to Problem 2.
A.4 Proposition 3.2
@@ B.3 Adaptive prediction sets
Appendix C Additional figures
Figure 9: A grid of our model’s outputs given randomly rotated, non-cherrypicked meshes from Shapenet. We render each mesh whose chamfer distance with respect to the ground truth is below a predetermined success threshold in blue, and we render the remainder in red. Our success threshold is the average chamfer distance between a shape randomly rotated by $10^{\circ}$ and the original shape; we compute this average across 100 Shapenet validation meshes. Even when predicted shapes do not agree with the ground truth orientation determined by the creators of Shapenet, their orientation is often acceptable in practice.
Figure 10: A grid of 20 adaptive prediction sets generated using our pipeline given randomly-rotated, non-cherrypicked meshes from Shapenet. We render each mesh whose chamfer distance with respect to the ground truth is below a predetermined success threshold in blue, and we render the remainder in red. Our success threshold is the average chamfer distance between a shape randomly rotated by $10^{\circ}$ and the original shape; we compute this average across 100 Shapenet validation meshes. Our adaptive prediction sets are typically small and include reasonable orientations, which an end user may choose between as needed for their target application.
Figure 11: A grid of our model’s outputs given randomly rotated, non-cherrypicked meshes from Objaverse. Our quotient regressor consistently succeeds on out of distribution meshes, as most of our pipeline’s outputs are correctly oriented up to a cube flip. Our flipper has greater difficulty generalizing, but predicts an acceptable flip in many cases. We expect that training on a larger and more diverse dataset of oriented shapes will improve our flipper’s generalization performance.